OpenStudy (anonymous):

Can someone please help me? Given ƒ(x) = x2 + 4 and g(x) = 5x − 3, find (ƒ ○ ƒ)(x) I got (fof)(x) = 5x^2+20 but i'm not sure it's the right answer

OpenStudy (anonymous):

@geerky42 do I do x^2(x^2+4)4 for this one?

OpenStudy (anonymous):

Given ƒ(x) = x^2 + 4 and g(x) = 5x − 3, find (ƒ ○ ƒ)(x)

geerky42 (geerky42):

No. You plug \(x^2+4\) into x. So you should have \((f\circ f)(x) = f(~f(x)~) = (x^2+4)^2+4\)

geerky42 (geerky42):